OpenStudy (anonymous):

Can someone explain to me how to solve this? Which of the following points lie in the solution set to the following system of inequalities? y ≤ x - 5 y ≥ -x - 4 A. (-5, 2) B. (5, -2) C. (-5, -2) D. (5, 2)

OpenStudy (agreene):

plug and chug would be the easiest way. the points are given in: (x,y) form, so you can plug them in for the equations, only one set will work on both

OpenStudy (anonymous):

I was thinking B

OpenStudy (agreene):

For the case of B): (5,-2) eqn 1: 2 ≤ 5-5 2 ≤ 0 <--- not true, can't be B. For the case of C: (-5,-2) eqn 1: -2 ≤ -5 - 5 -2 ≤ -10 <--- not true, can't be C As it happens, this also rules out A and D. the actual roots for this system are: x= 1/2; y =-9/2 and x>1/2; -x-4≤y≤x-5

OpenStudy (anonymous):

Wait? So, whats the answer? You ruled all of the answers out?

OpenStudy (anonymous):

and you made a mistake on the first equation, its suppose to be a negative 2, not positive.

OpenStudy (anonymous):

how am I suppose to answer it if all of the answer choices are wrong? this is confusing? lol

OpenStudy (agreene):

indeed i did, -2 ≤ 0 is true, y ≥ -x - 4 -2 ≥ -5 -4 -2 ≥ -9 is also true, I'm dumb, sorry... B is correct.

OpenStudy (anonymous):

haha its okay, I do it all the time, hahaha :D Thanks for your help! :)